OpenStudy (anonymous):

Solve the absolute value inequality: |x - 6| > 2 please help

OpenStudy (anonymous):

and just a quick answer please

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

i believe its |x| > -4

OpenStudy (artfis):

Change it to x - 6 = 2 and x- 6 = -2. Solve them both for the answer.

OpenStudy (anonymous):

no @SavageScopesFTW \[\left| x-6 \right|>2\] Get rid of the absolute sign. \[x-6>2\] or \[x-6<-2\]

OpenStudy (anonymous):

No @ArTFis You don't change it to an equal sign.

OpenStudy (anonymous):

thats not possible

OpenStudy (anonymous):

these are the possible answers: x > 4 or x < 8 x < 4 or x > 8 x < 4 and x > 8 x > 4 and x < 8

OpenStudy (anonymous):

@BigD;) I already gave you the setout.

OpenStudy (anonymous):

@Azteck well Arent you supposed to do the absolute value of the left section first then subtract both sides?